### v2

@@ Abstract
The complexity in finding a satisfactory explanation for the enhancement of in Fe alloys reveals a rich behavior in samples prepared under different procedures that can not be replicated for thin films. Interestingly, in bulk samples, the density of those nano inclusions involves annealing at high temperatures (1000 ∘C), quenching and aging at intermediate temperatures, and a second quenching [25]. These procedures are not accessible for thin film preparation aiming for flat topography and sharp interfaces since inter-diffusion with the substrate or buffer layers or the formation of rough layers with dome-like topography is activated by substrate temperature [26]. In any case, in the nanometric range, FeGa crystalline films with low magnetic anisotropy show an ME stress large enough to activate uniaxial magnetic anisotropies [4, 5] by piezostrains and/or ferrolastic strain, with large converse magnetoelectric coupling coefficients. The magneto-crystalline barrier for moving M in the (001) plane between easy and hard direction is (/4), which is below 10 kJm-3 for x18 [3] and 1 kJm-3 for x25 [27]. It can be overcome if uniaxial strain as low as 0.001 is applied, considering that can take values around 10 MPa, which is a value not unusual for thin films [5, 28] and about 4 times the Fe bulk value [29].
@@ 2.2 Characterization techniques
XRD diffraction was performed in the Bragg-Brentano configuration with a RIGAKU Ru2500 diffractometer (copper anode 40 kV/80 mA, with = 1.54 Å) and synchrotron radiation (BM25 Spain Beamline at ESRF, with = 0.62 Å). X-ray reflectometry was used to obtain the film thickness by fitting the intensity oscillations using a Bruker D8 diffractometer, Fig.1c. X-ray Absorption Fine Structure (XAFS) spectroscopy is a short-range-order, chemically selective technique complementary to diffraction. It can provide information on the atomic local environment of the different atoms comprising the samples (see ref [40] for an overview). The XAFS spectra were recorded with the beam polarization nearly parallel (, incidence angle 5∘) and almost perpendicular (, incidence angle 85∘) to the sample surface. We measured the extended region above the absorption edge of the XAFS spectrum (EXAFS) and the Near-Edge region (XANES) at the Fe and Ga K-edges at beamline I18 of the Diamond UK’s National Light Source. The raw polarized XANES spectra with parallel and perpendicular polarization were subtracted from each other to obtain the Linear Dichroism XANES signal (LDXANES) [41].
@@ 3.1 X-ray diffraction characterization
X-ray diffraction employing synchrotron radiation was used to look for SL reflections characteristic of D03 and mD03 structures, see Figure 2, in the 26-52 film, with the (001) reflection observed in the Brag-Brentano configuration. Here, we use the notation h, k, and l for the Miller indices for the A2 reciprocal space lattice, space group , Z = 2 and unit vector 2/, with the lattice parameter to be obtained but around 2.9 Å. The B2 structure, space group implies an order between Fe and Ga in the cell and many Ga-Ga pairs. For the D03 and mD03, the cell unit is doubled in real space, see Fig.1a, and ordering appears with fractional values of h, k, and l [43]. Using the D03 cell avoids the fractional notation and is straightforwardly done by multiplying by two the Miller indices. Figure 2.a shows a sketch of the reciprocal space with the reflections due to the A2 phase (h+k+l = even, black points) and SL reflections. For B2 crosses are added at h+k+l = odd; for the D phase distinguished reflections are added at [1/2](h k l), the red points, while for mD03 those specific SL peaks correspond to the green points at (h+0.5 k+0.5 l) for pairs along the [001], and (h+0.5 k l+0.5) or (h k+0.5 l+0.5) for in-plain [010] and [100] Ga pairs, see orange points. The lines indicate scans performed with synchrotron radiation to ascertain the presence of the D03 or mD03 phases and presented in Fig2 a and b. Those scans use the MgO crystal to set the reference frame (with H, K, and L Miller indeces and 2/ modulus of the unit cubic vector). The epitaxy between the MgO [110] parallel to the FeGa [100] of the bcc structure is observed at the scans with L , where peak overlapping is manifested, see Fig.2 b and c. Therefore, for scans along [HH.], h H, while for [H0.] h=k 2H. The more intense peaks correspond to the Bragg condition for the bcc lattice (h+k+l = even). Thus, [HH.] is parallel to [h0.], and the value of L for the FeGa reflection (l=1,2, etc.) is determined by maximizing the intensity for the FeGa (002) reflection. For l=1, blue lines, the A2 (101), (301), and (501) reflections, and the SL peak (111) are visible.
@@ 6 Acknowledgments
PID2021-124734OB-C21 fun-ded by MCIN /AEI /10.13039/501100011033 and by “ERDF A way of making Europe” and grant CEX2023-001286-S funded by MICIU/AEI /10.13039/501100011033 and Gobierno de Aragón (grant E12-23R). A. B. would like to acknowledge the funding received from the Ministry of Universities and the European Union-Next Generation for the Margarita Salas fellowship. We are grateful to the staff of beamline I18 at Diamond UK’s National Synchrotron Light source and BM25 Spain Beamline at ESRF for their strong support in performing the experiments.

### v3

@@ Abstract
The complexity in finding a satisfactory explanation for the enhancement of in Fe alloys reveals a rich behavior in samples prepared under different procedures that can not be replicated for thin films. Interestingly, in bulk samples, the density of those nano inclusions involves annealing at high temperatures (1000 ∘C), quenching and aging at intermediate temperatures, and a second quenching [25]. These procedures are not accessible for thin film preparation aiming for flat topography and sharp interfaces since inter-diffusion with the substrate or buffer layers or the formation of rough layers with dome-like topography is activated by substrate temperature [26]. In any case, in the nanometric range, FeGa crystalline films with low magnetic anisotropy show an ME stress large enough to activate uniaxial magnetic anisotropies [4, 5] by piezostrains and/or ferroelastic strain, with large converse magnetoelectric coupling coefficients. The magneto-crystalline barrier for moving M in the (001) plane between easy and hard direction is (/4), which is below 10 kJm-3 for x18 [3] and 1 kJm-3 for x25 [27]. It can be overcome if uniaxial strain as low as 0.001 is applied, considering that can take values around 10 MPa, which is a value not unusual for thin films [5, 28] and about 4 times the Fe bulk value [29].
@@ 2.2 Characterization techniques
XRD diffraction was performed in the Bragg-Brentano configuration with a RIGAKU Ru2500 diffractometer (copper anode 40 kV/80 mA, with = 1.54 Å) and synchrotron radiation (BM25 Spain Beamline at ESRF, with = 0.62 Å). X-ray reflectometry was used to obtain the film thickness by fitting the intensity oscillations using a Bruker D8 diffractometer, Fig.1c. X-ray Absorption Fine Structure (XAFS) spectroscopy is a short-range-order, chemically selective technique complementary to diffraction. It can provide information on the atomic local environment of the different atoms comprising the samples (see ref [40] for an overview). The XAFS spectra were recorded with the beam polarization nearly parallel (, incidence angle 5∘) and almost perpendicular (, incidence angle 85∘) to the sample surface. We measured the extended region above the absorption edge of the XAFS spectrum (EXAFS) and the Near-Edge region (XANES) at the Fe and Ga K-edges at beamline B18 of the Diamond UK’s National Light Source. The raw polarized XANES spectra with parallel and perpendicular polarization were subtracted from each other to obtain the Linear Dichroism XANES signal (LDXANES) [41].
@@ 3.1 X-ray diffraction characterization
X-ray diffraction employing synchrotron radiation was used to look for SL reflections characteristic of D03 and mD03 structures, see Figure 2, in the 26-52 film, with the (001) reflection observed in the Brag-Brentano configuration. Here, we use the notation h, k, and l for the Miller indices for the A2 reciprocal space lattice, space group , Z = 2 and unit vector 2/, with the lattice parameter to be obtained but around 2.9 Å. The B2 structure, space group implies an order between Fe and Ga in the cell and many Ga-Ga pairs. For the D03 and mD03, the cell unit is doubled in real space, see Fig.1a, and ordering appears with fractional values of h, k, and l [43]. Using the D03 cell avoids the fractional notation and is straightforwardly done by multiplying by two the Miller indices. Figure 2.a shows a sketch of the reciprocal space with the reflections due to the A2 phase (h+k+l = even, black points) and SL reflections. For B2 crosses are added at h+k+l = odd; for the D phase distinguished reflections are added at [1/2](h k l), the red points, while for mD03 those specific SL peaks correspond to the green points at (h+0.5 k+0.5 l) for pairs along the [001], and (h+0.5 k l+0.5) or (h k+0.5 l+0.5) for in-plain [010] and [100] Ga pairs, see orange points. The lines indicate scans performed with synchrotron radiation to ascertain the presence of the D03 or mD03 phases and presented in Fig2 a and b. Those scans use the MgO crystal to set the reference frame (with H, K, and L Miller indices and 2/ modulus of the unit cubic vector). The epitaxy between the MgO [110] parallel to the FeGa [100] of the bcc structure is observed at the scans with L , where peak overlapping is manifested, see Fig.2 b and c. Therefore, for scans along [HH.], h H, while for [H0.] h=k 2H. The more intense peaks correspond to the Bragg condition for the bcc lattice (h+k+l = even). Thus, [HH.] is parallel to [h0.], and the value of L for the FeGa reflection (l=1,2, etc.) is determined by maximizing the intensity for the FeGa (002) reflection. For l=1, blue lines, the A2 (101), (301), and (501) reflections, and the SL peak (111) are visible.
@@ 6 Acknowledgments
PID2021-124734OB-C21 fun-ded by MCIN /AEI /10.13039/501100011033 and by “ERDF A way of making Europe” and grant CEX2023-001286-S funded by MICIU/AEI /10.13039/501100011033 and Gobierno de Aragón (grant E12-23R). A. B. would like to acknowledge the funding received from the Ministry of Universities and the European Union-Next Generation for the Margarita Salas fellowship. We are grateful to the staff of beamline B18 at Diamond UK’s National Synchrotron Light source and BM25 Spain Beamline at ESRF for their strong support in performing the experiments.